OpenStudy (anonymous):

Line CD passes through points C(1, 3) and D(4, –3). If the equation of the line is written in slope-intercept form, y = mx + b, what is the value of b? A)–5 B)–2 C)1 D)5

OpenStudy (mathmate):

hints: 1. find m (slope) = (y2-y1)/(x2-x1) 2. substitute one of the two points in the equation y=mx+b to get b=y-mx, or b=y1-m(x1) if point #1 is used.